OpenStudy (anonymous):

Hey, could anyone help me?? Square rooted 29+7+3x2 divide by 3

OpenStudy (anonymous):

\[\sqrt{29+7}+ 3^{2} \div3\]

OpenStudy (anonymous):

9

OpenStudy (anonymous):

\[\sqrt{36}+3^{2-1}= 6+3=9\]

OpenStudy (anonymous):

can you understand?

OpenStudy (anonymous):

Yea! thank you both!!

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

Who do I give the medal to??

OpenStudy (anonymous):

it's more important that you surely understand